# Physics-Inspired Modeling and Content Adaptive Routing in an Infrared Gas Leak Detection Network

### v1

@@ Abstract
Detecting infrared gas leaks is critical for environmental monitoring and industrial safety, yet remains difficult because plumes are faint, small, semitransparent, and have weak, diffuse boundaries. We present physics-edge hybrid gas dynamic routing network (PEG-DRNet). First, we introduce the Gas Block, a diffusion-convection unit modeling gas transport: a local branch captures short-range variations, while a large-kernel branch captures long-range propagation. An edge-gated learnable fusion module balances local detail and global context, strengthening weak-contrast plume and contour cues. Second, we propose the adaptive gradient and phase edge operator (AGPEO), computing reliable edge priors from multi-directional gradients and phase-consistent responses. These are transformed by a multi-scale edge perception module (MSEPM) into hierarchical edge features that reinforce boundaries. Finally, the content-adaptive sparse routing path aggregation network (CASR-PAN), with adaptive information modulation modules for fusion and self, selectively propagates informative features across scales based on edge and content cues, improving cross-scale discriminability while reducing redundancy. Experiments on the IIG dataset show that PEG-DRNet achieves an overall AP of 29.8%, an AP50 of 84.3%, and a small-object AP of 25.3%, surpassing the RT-DETR-R18 baseline by 3.0%, 6.5%, and 5.3%, respectively, while requiring only 43.7 Gflops and 14.9 M parameters. The proposed PEG-DRNet achieves superior overall performance with the best balance of accuracy and computational efficiency, outperforming existing CNN and Transformer detectors in AP and AP50 on the IIG and LangGas dataset.
@@
Infrared gas leak detection, Deep learning, Physics-Inspired modeling, Content-adaptive sparse route path aggregation network.
@@
Physics-edge hybrid gas dynamic routing network.
Novel edge detection operator based on adaptive gradient and phase consistency.
Content-adaptive sparse route path aggregation network for dynamic multi-scale feature fusion with selective path aggregation.
@@ 1 Introduction
With the rapid development of the petrochemical and natural gas industries, these sectors have provided strong support for industrialization and urbanization processes (Xu et al., 2023). However, the dense layout of industrial facilities and the extensive scale of pipeline networks have also made gas leak accidents a major concern for industrial safety and environmental protection (Lu et al., 2020). Substantial experimental evidence and accident reports indicate that gas leaks can not only directly trigger catastrophic events such as fires and explosions, but can also lead to large-scale release of toxic and hazardous substances in the early stages of leakage, thus exerting profound and long-lasting impacts on ecosystems and public health (Zuo et al., 2023; Bonvicini et al., 2015; Kopbayev et al., 2022). Therefore, achieving efficient gas leak detection has become an urgent problem that needs to be addressed.
@@
In contrast, infrared (IR) thermography does not rely on physical contact and enables remote capture of temperature distributions in target areas, which is particularly valuable for online monitoring in hazardous environments (Strahl et al., 2021). IR imaging can be classified into active and passive modalities. Active imaging requires an external light source, offering high sensitivity, but its safety and detectability are constrained by the external source. In gas leakage scenarios, temperature and pressure typically decrease, creating thermal differences between the leaked gas and the background, which satisfies the conditions for passive detection. Many hazardous industrial gases exhibit characteristic absorption spectra in the 3-17 µm band. When dynamic leakage occurs within the field of view of an IR camera, radiation differences arise at the corresponding absorption wavelengths, generating grayscale contrast in the captured images. The higher the gas concentration, the stronger the absorption and the more pronounced the grayscale contrast. Furthermore, as gas plumes exhibit drifting states, plume detection and concentration estimation can be achieved through analysis of concentration differences and dynamic motion.
With the rapid advancement of deep learning, particularly convolutional neural networks (CNNs) and Transformer-based architectures, IR gas leak detection has embraced new opportunities. Deep learning has been widely applied in image processing and object detection, where its powerful capability for automated feature extraction and representation has driven significant breakthroughs beyond traditional methods (Li et al., 2025; Zhang et al., 2025; Zhong et al., 2024; Wang et al., 2025). However, applying deep learning to IR gas leak detection still faces two critical challenges. First, gas diffusion is inherently dynamic and unstable, often exhibiting turbulence and plume-like patterns; its concentration decreases gradually from the center outward, resulting in indistinct boundaries. Moreover, thermal interactions with the environment lead to low-contrast images, while partial gas radiation overlaps with the background, making features weak and difficult to recognize accurately. Second, existing models often suffer from insufficient feature extraction capacity and high computational cost.
Based on this, a single-stage detection model with the following contributions is proposed.It strikes a good balance between high accuracy and computational efficiency. Figure. 1 shows the comparison results of the proposed method with leading methods in terms of accuracy and computational complexity on the IIG dataset.
@@ 2.2 Infrared Gas Leak Detection Based on Deep Learning
Deep learning-based infrared gas leak detection is typically built upon generic object detection frameworks, which can be categorized into two-stage and one-stage paradigms. Two-stage detectors, such as R-CNN (Girshick et al., 2014), Fast R-CNN (Girshick, 2015), and Faster R-CNN (Ren et al., 2016), employ region proposals followed by classification and bounding-box refinement, achieving high accuracy at the cost of slower inference. For example, TSFF-Net (Yao et al., 2024) fuses temporal motion and spatial details to improve SF6 leak detection in complex scenes, but its reliance on motion cues and sequential computation makes static leak detection and edge deployment challenging.
@@
Beyond object detectors, edge detection and feature pyramid architectures are widely adopted to enhance small or low-contrast target representation. Classical operators such as Sobel, Canny, and Laplacian (Heath et al., 1998; Canny, 2009; Wang, 2007) emphasize intensity variations and capture structural cues, but are highly sensitive to noise, perform poorly under low contrast or semi-transparent conditions, and lack semantic awareness. Modern detectors thus integrate multi-scale feature pyramids, including FPN (Lin et al., 2017), PANet (Liu et al., 2018), BiFPN (Tan et al., 2020), and NAS-FPN (Ghiasi et al., 2019), to fuse high-level semantics with low-level details. However, their fusion strategies are typically static and content-agnostic, limiting adaptability to weak-edge or content-varying gas plumes.
In contrast to the above, we propose a physics-edge hybrid gas dynamic routing network (PEG-DRNet). The Gas Block integrates local diffusion and global convection to enhance weak-contrast plume and contour representations, while the multi-scale edge perception module (MSEPM), built upon an adaptive gradient and phase edge operator (AGPEO), produces reliable edge priors under semi-transparent, low-contrast conditions. Furthermore, the content-adaptive sparse routing path aggregation network (CASR-PAN) dynamically modulates cross-scale information flow via adaptive information modulation modules for fusion(AIMM-F) and adaptive information modulation modules for self(AIMM-S), selectively propagating informative features and suppressing redundancy. This design bridges classical physics-based modeling and modern deep architectures, improving robust detection of faint and semi-transparent gas plumes under challenging industrial conditions.
@@ 3.3 Multi-Scale Edge Perception Module
To address this issue, the multi-scale edge perception module (MSEPM), which combines multi-directional gradient detection with phase consistency analysis to preserve subtle gas edge information while effectively suppressing background noise, is introduced. Specifically, the module first employs the adaptive gradient and phase edge operator (AGPEO) to adaptively fuse gradient and phase information, enhancing the edge representation of gas targets. Subsequently, multi-scale pooling and feature transformation are applied to generate edge feature maps corresponding to the shallow, middle, and deep layers of the backbone network, providing more discriminative edge representations for subsequent feature fusion and target detection. The module architecture is illustrated in Figure. 4.
@@
where, $E_{0}$ denotes the initial edge feature map, which retains weak gas boundaries while suppressing background noise.The fusion weight $\alpha$ is implemented as a learnable scalar parameter and optimized jointly with the network via back-propagation under the detection loss.This design enables adaptive balancing between gradient-based and phase-based cues in a data-driven manner, without introducing additional supervision or manual tuning.
@@ 3.4 Content-Adaptive Sparse Route Path Aggregation Network
In contrast, the proposed content-adaptive sparse routing path aggregation network (CASR-PAN) performs fully content-driven multi-scale fusion. As illustrated in Figure. 1, CASR-PAN dynamically generates spatially varying path weights through an importance estimator (IE), selectively enhancing or suppressing cross-scale feature flows using sparse routing. Multi-path outputs at the same scale are further consolidated by RepC3 modules, achieving stable gradient propagation and discriminative multi-scale representation. The internal structure of the IE is shown in Figure. 5.
@@
The three components $\tilde{\mathbf{G}}$, $\mathbf{L}$, and $\mathbf{D}$ are designed to capture complementary aspects of feature importance. Specifically, $\tilde{\mathbf{G}}$ represents the global importance. It is obtained by applying global average pooling to $\mathbf{X}$, followed by a $1\times 1$ convolution, ReLU activation, another $1\times 1$ convolution with sigmoid, and upsampling to the original spatial resolution. The global average pooling aggregates information across all spatial locations, producing a vector that represents the overall contribution of each channel to the feature map, independent of spatial position. The subsequent $1\times 1$ convolutions allow non-linear interactions between channels, enhancing the representation of channels that are globally significant. This combination effectively captures holistic, channel-wise importance, which is why it serves as a “global” measure: it summarizes each channel’s overall influence rather than local variations. The local importance $\mathbf{L}$ is computed using a $3\times 3$ convolution followed by ReLU, a $1\times 1$ convolution, and sigmoid activation. A $3\times 3$ convolution operates on a small neighborhood of each spatial location, naturally focusing on local spatial dependencies. By scanning the feature map with a limited receptive field, it captures fine-grained variations, such as edges, textures, or localized patterns, that may not be reflected in global pooling. This design ensures that the estimator can adaptively weigh spatially local features, distinguishing important regions from surrounding areas. The following $1\times 1$ convolution allows channel-wise integration while maintaining spatial resolution. The diversity component $\mathbf{D}$ estimates feature diversity via $1\times 1$ convolutions followed by ReLU and another $1\times 1$ convolution with sigmoid, further modulated by the per-channel standard deviation of $\mathbf{X}$. Where $1\times 1$ convolutions operate independently at each spatial location, focusing solely on cross-channel interactions. Combined with the statistical measure of variation, this highlights channels that carry complementary or non-redundant information. The use of $1\times 1$ convolutions ensures that diversity is measured across channels rather than across space, emphasizing the variety and richness of feature representations rather than spatial structure. By combining these global, local, and diversity cues, with carefully chosen operations that align with their conceptual goals, the IE effectively produces a content-adaptive importance map that guides the sparse route path aggregation in CASR-PAN, enabling the network to prioritize informative features while suppressing less relevant ones.
@@
    $F_{\text{local}}(x,y)$: the feature from the current scale,
@@ 4.1 Dataset
The IIG dataset was collected by Yu et al. (Yu et al., 2024)at Zhejiang University using a handheld infrared camera (VF 330-1000 thermal infrared imager), containing a total of 11,186 instances of gas leakage. The dataset is split into a training set and a test set in a 4:1 ratio, with the training set containing 4,453 images and 9,289 gas leakage instances, and the test set containing 1,116 images and 1,897 gas leakage instances. To further verify the effectiveness and generalization ability of the proposed method, we additionally employed the LangGas dataset (Guo et al., 2025) in our experiments. LangGas is a computer-synthetic infrared gas leakage dataset designed to advance research in this field. It provides diverse leakage patterns, perfectly accurate pixel-level segmentation ground truth, and stable video sequences containing multiple moving foreground objects. The dataset is generated by overlaying realistically rendered gas plumes and various interfering objects onto a wide range of background scenes. For our study, we manually annotated images for object detection purposes and divided the data into a training set and a validation set. The training set contains 1289 images, and the validation set contains 323 images.
@@ 4.3 Comparative Experiments
To evaluate the effectiveness of the proposed PEG-DRNet, several representative object detection algorithms, including YOLOv10 (Wang et al., 2024), YOLOv11 (Khanam and Hussain, 2024), YOLOv12 (Tian et al., 2025), SSD (Liu et al., 2016), Faster R-CNN (Ren et al., 2016), and RT-DETR (Zhao et al., 2024), were selected as baselines for comparison. The quantitative results are summarized in Table  1, highlighting the performance advantages of PEG-DRNet over these existing methods.
@@
Model AP (%) AP50 (%) AP75 (%) APS (%) APM (%) APL (%) Gflops (G) Parameters (M)
@@
Model AP (%) AP50 (%) AP75 (%) APS (%) APM (%) APL (%) Gflops (G) Parameters (M)
@@
Our model achieves the highest AP of 29.8%, outperforming all baselines (e.g., +3.0% over RT-DETR-R18 and +3.6% over Yolov10n), which demonstrates the benefits of integrating Physics-Inspired modeling, edge-aware representation, and adaptive multiscale fusion. Notably, PEG-DRNet attains a substantial improvement in small-object detection (APS=25.3%, +5.3% over Yolov12n), indicating that fine-grained spatial details are well preserved, while medium- and large-object detection (APM=32.5%, APL=42.6%) remains competitive, highlighting balanced scale adaptability. Despite superior accuracy, the model requires only 43.7 Gflops and 14.932M parameters, lower than heavier baselines such as Faster-RCNN (134.5 Gflops, 41.35M), demonstrating practical deployability. The high AP50=84.3% and reasonable AP75=8.5% suggest effective localization even for subtle leaks, though the AP75 gap implies that extremely faint or overlapping leaks remain challenging, likely due to sensor noise or cluttered backgrounds.
@@
To further validate the effectiveness of the proposed method, a visualization comparison experiment was conducted. The YOLO series employed the YOLOv12n, which currently achieves the best detection performance among its variants. The comparative visualization results are presented in Figure. 6.Although YOLOv12 has relatively fewer parameters and lower computational complexity, the visualization comparison reveals that it suffers from a high miss detection rate, which is a critical issue in gas leak detection tasks. The single-stage detector SSD exhibits similar shortcomings, indicating its limited sensitivity to faint and small gas plumes. In contrast, Faster R-CNN achieves higher detection accuracy than the previous two models; however, its heavy model parameters and substantial computational burden hinder deployment on resource-constrained edge devices. The RT-DETR-R18 model offers a better trade-off between detection accuracy and computational efficiency, yet it still exhibits false positives and missed detections under challenging conditions.In comparison, the proposed PEG-DRNet demonstrates superior localization capability and successfully identifies most gas plumes, outperforming the mainstream detection models. Nevertheless, it still struggles with extremely small and visually indistinct gas leaks at long distances. Through comprehensive analysis of these scenarios, it is concluded that gas detection accuracy decreases in visually complex and cluttered backgrounds compared with simple and uniform scenes. Moreover, scenes containing larger and more structured gas plumes tend to yield higher detection accuracy. Future research will focus on enhancing the detection precision of subtle and blurry gas leaks under complex backgrounds, which is expected to further advance the development of vision-based infrared gas detection systems.

### v2

@@ Abstract
Detecting infrared gas leaks is critical for environmental monitoring and industrial safety, yet remains difficult because plumes are faint, small, semitransparent, and have weak and diffuse boundaries. We present a physics-edge hybrid gas dynamic routing network (PEG-DRNet). First, we introduce the Gas Block, a diffusion-convection unit modeling gas transport: a local branch captures short-range variations, while a global (large-kernel) branch captures long-range propagation. An edge-gated learnable fusion module balances local detail and global context, strengthening weak-contrast plume and contour cues. Second, we propose the adaptive gradient and phase edge operator (AGPEO), computing reliable edge priors from multi-directional gradients and phase-consistent responses. These are transformed by a multi-scale edge perception module (MSEPM) into hierarchical edge features that reinforce boundaries. Finally, we propose the content-adaptive sparse routing path aggregation network (CASR-PAN). The network uses adaptive information modulation modules for fusion and self. It selectively propagates informative features across scales according to edge and content cues, enhancing cross-scale discriminability and reducing redundancy. Experiments on the IIG dataset show that PEG-DRNet achieves an overall AP of 29.8%, an AP50 of 84.3%, and a small-object AP of 25.3%, surpassing the RT-DETR-R18 baseline by 3.0%, 6.5%, and 5.3%, respectively, while requiring only 43.7 Gflops and 14.9 M parameters. The proposed PEG-DRNet achieves superior overall performance with the best balance of accuracy and computational efficiency, outperforming existing CNN and Transformer detectors in AP and AP50 on the IIG and LangGas dataset.
@@
Infrared gas leak detection, Deep learning, Physics-Inspired modeling, CASR-PAN.
@@
[…]
[3] organization=Shaanxi Shanhua Coal Chemical Co.,Ltd, city=Weinan, postcode=714104, state=Shaanxi, country=China
@@
A physics-edge hybrid gas dynamic routing network.
A novel edge detection operator based on adaptive gradient and phase consistency.
A content-adaptive sparse routing path aggregation network for dynamic multi-scale feature fusion.
@@ 1 Introduction
With the rapid development of the petrochemical and natural gas industries, these sectors have provided strong support for industrialization and urbanization processes (Xu et al., 2023). However, the dense layout of industrial facilities and the extensive scale of pipeline networks have also made gas leak accidents a major concern for industrial safety and environmental protection (Lu et al., 2020). Substantial experimental evidence and accident reports indicate that gas leaks can not only directly trigger catastrophic events such as fires and explosions, but can also lead to large-scale release of toxic and hazardous substances in the early stages of leakage, thus exerting profound and long-lasting impacts on ecosystems and public health (Zuo et al., 2024; Bonvicini et al., 2015; Kopbayev et al., 2022). Therefore, achieving efficient gas leak detection has become an urgent problem that needs to be addressed.
@@
In contrast, infrared (IR) thermography enables non-contact measurement, allowing temperature distributions to be captured remotely. This feature is particularly advantageous for online monitoring in hazardous environments (Strahl et al., 2021). IR imaging can be classified into active and passive modalities. Active imaging requires an external light source, offering high sensitivity, but its safety and detectability are constrained by the external source. In gas leakage scenarios, temperature and pressure typically decrease, creating thermal differences between the leaked gas and the background, which satisfies the conditions for passive detection. Many hazardous industrial gases exhibit characteristic absorption spectra in the 3-17 µm band. When dynamic leakage occurs within the field of view of an IR camera, radiation differences arise at the corresponding absorption wavelengths, generating grayscale contrast in the captured images. The higher the gas concentration, the stronger the absorption and the more pronounced the grayscale contrast. Furthermore, as gas plumes exhibit drifting states, plume detection and concentration estimation can be achieved through analysis of concentration differences and dynamic motion.
With the rapid advancement of deep learning, particularly convolutional neural networks (CNNs) and Transformer-based architectures, IR gas leak detection has embraced new opportunities. Deep learning has been widely applied in image processing and object detection, where its powerful capability for automated feature extraction and representation has driven significant breakthroughs beyond traditional methods (Li et al., 2026; Zhang et al., 2025; Zhong et al., 2024; Wang et al., 2026). However, applying deep learning to IR gas leak detection still faces two critical challenges. First, gas diffusion is inherently dynamic and unstable, often exhibiting turbulence and plume-like patterns; its concentration decreases gradually from the center outward, resulting in indistinct boundaries. Moreover, thermal interactions with the environment lead to low-contrast images, while partial gas radiation overlaps with the background, making features weak and difficult to recognize accurately. Second, existing models often suffer from insufficient feature extraction capacity and high computational cost.
To address these challenges, a single-stage detection model with the following contributions is proposed.It strikes a good balance between high accuracy and computational efficiency. Figure 1 shows the comparison results of the proposed method with leading methods in terms of accuracy and computational complexity on the IIG dataset.
@@ 2.2 Infrared Gas Leak Detection Based on Deep Learning
Deep learning-based infrared gas leak detection is typically built upon generic object detection frameworks, which can be categorized into two-stage and one-stage paradigms. Two-stage detectors, such as R-CNN (Girshick et al., 2014), Fast R-CNN (Girshick, 2015), and Faster R-CNN (Ren et al.,, 2017), employ region proposals followed by classification and bounding-box refinement, achieving high accuracy at the cost of slower inference. For example, TSFF-Net (Yao et al., 2024) fuses temporal motion and spatial details to improve SF6 leak detection in complex scenes, but its reliance on motion cues and sequential computation makes static leak detection and edge deployment challenging.
@@
Beyond object detectors, edge detection and feature pyramid architectures are widely adopted to enhance small or low-contrast target representation. Classical operators such as Sobel, Canny, and Laplacian (Heath et al., 1998; Canny, 1986; Wang, 2007) emphasize intensity variations and capture structural cues, but are highly sensitive to noise, perform poorly under low contrast or semi-transparent conditions, and lack semantic awareness. Modern detectors thus integrate multi-scale feature pyramids, including FPN (Lin et al., 2017), PANet (Liu et al., 2018), BiFPN (Tan et al., 2020), and NAS-FPN (Ghiasi et al., 2019), to fuse high-level semantics with low-level details. However, their fusion strategies are typically static and content-agnostic, limiting adaptability to weak-edge or content-varying gas plumes.
To address the aforementioned limitations, we propose a physics-edge hybrid gas dynamic routing network (PEG-DRNet). The Gas Block integrates local diffusion and global convection to enhance weak-contrast plume and contour representations, while the multi-scale edge perception module (MSEPM), built upon an adaptive gradient and phase edge operator (AGPEO), produces reliable edge priors under semi-transparent, low-contrast conditions. Furthermore, the content-adaptive sparse routing path aggregation network (CASR-PAN) dynamically modulates cross-scale information flow via adaptive information modulation modules for fusion(AIMM-F) and adaptive information modulation modules for self(AIMM-S), selectively propagating informative features and suppressing redundancy. This design bridges classical physics-based modeling and modern deep architectures, improving robust detection of faint and semi-transparent gas plumes under challenging industrial conditions.
@@ 3.3 Multi-Scale Edge Perception Module
To address this issue, the multi-scale edge perception module (MSEPM), which combines multi-directional gradient detection with phase consistency analysis to preserve subtle gas edge information while effectively suppressing background noise, is introduced. Specifically, the module first employs the adaptive gradient and phase edge operator (AGPEO) to adaptively fuse gradient and phase information, enhancing the edge representation of gas targets. Subsequently, multi-scale pooling and feature transformation are applied to generate edge feature maps corresponding to the shallow, middle, and deep layers of the backbone network, providing more discriminative edge representations for subsequent feature fusion and target detection. The module architecture is illustrated in Figure 4.
@@
where $E_{0}$ denotes the initial edge feature map, which retains weak gas boundaries while suppressing background noise.The fusion weight $\alpha$ is implemented as a learnable scalar parameter and optimized jointly with the network via back-propagation under the detection loss.This design enables adaptive balancing between gradient-based and phase-based cues in a data-driven manner, without introducing additional supervision or manual tuning.
@@ 3.4 Content-Adaptive Sparse Route Path Aggregation Network
In contrast, the proposed content-adaptive sparse routing path aggregation network (CASR-PAN) performs fully content-driven multi-scale fusion. As illustrated in Figure 2, CASR-PAN dynamically generates spatially varying path weights through an importance estimator (IE), selectively enhancing or suppressing cross-scale feature flows using sparse routing. Multi-path outputs at the same scale are further consolidated by RepC3 modules, achieving stable gradient propagation and discriminative multi-scale representation. The internal structure of the IE is shown in Figure 5.
@@
The three components $\tilde{\mathbf{G}}$, $\mathbf{L}$, and $\mathbf{D}$ are designed to capture complementary aspects of feature importance. Specifically, $\tilde{\mathbf{G}}$ represents the global importance. It is obtained by applying global average pooling to $\mathbf{X}$, followed by a $1\times 1$ convolution, ReLU activation, another $1\times 1$ convolution with sigmoid, and upsampling to the original spatial resolution. The global average pooling aggregates information across all spatial locations, producing a vector that represents the overall contribution of each channel to the feature map, independent of spatial position. The subsequent $1\times 1$ convolutions allow non-linear interactions between channels, enhancing the representation of channels that are globally significant. This combination effectively captures holistic, channel-wise importance, which is why it serves as a “global” measure: it summarizes each channel’s overall influence rather than local variations. The local importance $\mathbf{L}$ is computed using a $3\times 3$ convolution followed by ReLU, a $1\times 1$ convolution, and sigmoid activation. A $3\times 3$ convolution operates on a small neighborhood of each spatial location, naturally focusing on local spatial dependencies. By scanning the feature map with a limited receptive field, it captures fine-grained variations, such as edges, textures, or localized patterns, that may not be reflected in global pooling. This design ensures that the estimator can adaptively weigh spatially local features, distinguishing important regions from surrounding areas. The following $1\times 1$ convolution allows channel-wise integration while maintaining spatial resolution. The diversity component $\mathbf{D}$ estimates feature diversity via $1\times 1$ convolutions followed by ReLU and another $1\times 1$ convolution with sigmoid, further modulated by the per-channel standard deviation of $\mathbf{X}$. Where, $1\times 1$ convolutions operate independently at each spatial location, focusing solely on cross-channel interactions. Combined with the statistical measure of variation, this highlights channels that carry complementary or non-redundant information. The use of $1\times 1$ convolutions ensures that diversity is measured across channels rather than across space, emphasizing the variety and richness of feature representations rather than spatial structure. By combining these global, local, and diversity cues, with carefully chosen operations that align with their conceptual goals, the IE effectively produces a content-adaptive importance map that guides the sparse route path aggregation in CASR-PAN, enabling the network to prioritize informative features while suppressing less relevant ones.
@@
    $F_{\text{local}}(x,y)$: the feature from the current scale.
@@ 4.1 Dataset
Experiments were carried out on two benchmark infrared gas leakage datasets. The IIG dataset, proposed by Yu et al. (Yu et al., 2024), was acquired at Zhejiang University using a VF 330-1000 handheld thermal infrared imaging device and contains a total of 11,186 annotated gas leakage instances. In accordance with the dataset protocol, the images were partitioned into training and testing subsets at a ratio of 4:1, resulting in 4,453 training images with 9,289 instances and 1,116 testing images with 1,897 instances. In addition, the LangGas dataset (Guo et al., 2025) was utilized to assess the generalization capability of the proposed approach. LangGas is a synthetically generated infrared gas leakage dataset constructed by compositing physically realistic gas plumes and various interfering foreground objects onto diverse background scenes. Although the dataset provides pixel-level segmentation annotations and video sequences, we manually labeled the images for object detection tasks. The annotated data were divided into 1,289 training images and 323 validation images.
@@ 4.3 Comparative Experiments
To evaluate the effectiveness of the proposed PEG-DRNet, several representative object detection algorithms, including YOLOv10 (Wang et al., 2024), YOLOv11 (Khanam and Hussain, 2024), YOLOv12 (Tian et al., 2025), SSD (Liu et al., 2016), Faster R-CNN (Ren et al.,, 2017), and RT-DETR (Zhao et al., 2024), were selected as baselines for comparison. The quantitative results are summarized in Table  1, highlighting the performance advantages of PEG-DRNet over these existing methods.
@@
Model AP (%) AP50 (%) AP75 (%) APS (%) APM (%) APL (%) Gflops (G) Params (M)
@@
Model AP (%) AP50 (%) AP75 (%) APS (%) APM (%) APL (%) Gflops (G) Params (M)
@@
Our model achieves the highest AP of 29.8%, outperforming all baselines (e.g., +3.0% over RT-DETR-R18 and +3.6% over Yolov10n), which demonstrates the benefits of integrating Physics-Inspired modeling, edge-aware representation, and adaptive multiscale fusion. Notably, PEG-DRNet attains a substantial improvement in small-object detection (APS=25.3%, +5.3% over Yolov12n), indicating that fine-grained spatial details are well preserved, while medium- and large-object detection (APM=32.5%, APL=42.6%) remains competitive, highlighting balanced scale adaptability. Despite superior accuracy, the model requires only 43.7 Gflops and 14.932M parameters, lower than heavier baselines such as Faster-RCNN (134.5 Gflops, 41.35M), demonstrating practical deployability. The high AP50{50}=84.3% confirms reliable region-level detection, whereas the comparatively lower AP75{75} reflects the intrinsic difficulty of precise bounding-box alignment for diffuse and semi-transparent gas plumes in infrared imagery, where ambiguous boundaries, and background clutter limit high-IoU localization; addressing such fine-grained localization remains a challenging direction for future work.
@@
To further validate the effectiveness of the proposed method, a visualization comparison experiment was conducted. The YOLO series employed the YOLOv12n, which currently achieves the best detection performance among its variants. The comparative visualization results are presented in Figure 6. Although YOLOv12 has relatively fewer parameters and lower computational complexity, the visualization comparison reveals that it suffers from a high false negative rate (FNR), which is a critical issue in gas leak detection tasks. The single-stage detector SSD exhibits similar shortcomings, indicating its limited sensitivity to faint and small gas plumes. In contrast, Faster R-CNN achieves higher detection accuracy than the previous two models; however, its heavy model parameters and substantial computational burden hinder deployment on resource-constrained edge devices. The RT-DETR-R18 model offers a better trade-off between detection accuracy and computational efficiency, yet it still exhibits false positives and missed detections under challenging conditions.In comparison, the proposed PEG-DRNet demonstrates superior localization capability and successfully identifies most gas plumes, outperforming the mainstream detection models. Nevertheless, it still struggles with extremely small and visually indistinct gas leaks at long distances. Through comprehensive analysis of these scenarios, it is concluded that gas detection accuracy decreases in visually complex and cluttered backgrounds compared with simple and uniform scenes. Moreover, scenes containing larger and more structured gas plumes tend to yield higher detection accuracy. Future research will focus on enhancing the detection precision of subtle and blurry gas leaks under complex backgrounds, which is expected to further advance the development of vision-based infrared gas detection systems.